# Anomalous Coherence Length in Superconductors with Quantum Metric

### v5

@@ V.1 Continuum model of twisted bilayer graphene
$H_{int}=\sum_{\xi}a^{\dagger}_{1,\xi,\bm{k}}[\left(\begin{array}[]{cc}w_{0}&w_{1}\\ w_{1}&w_{0}\\ \end{array}\right)\delta_{\bm{k},\bm{k}^{\prime}}+\left(\begin{array}[]{cc}w_{0}&w_{1}e^{-i\frac{2\pi}{3}}\\ w_{1}e^{i\frac{2\pi}{3}}&w_{0}\\ \end{array}\right)\delta_{\bm{k+\bm{G}_{1}^{M}},\bm{k}^{\prime}}+\left(\begin{array}[]{cc}w_{0}&w_{1}e^{i\frac{2\pi}{3}}\\ w_{1}e^{-i\frac{2\pi}{3}}&w_{0}\\ \end{array}\right)\delta_{\bm{k}+\bm{G}_{1}^{M}+\bm{G}_{2}^{M},\bm{k}^{\prime}}]a_{2,\xi,\bm{k}^{\prime}}+h.c.~{}.$ (S89)
Figure S3: The moiré Brillouin zone (mBZ) of moiré graphene family. The original Brillouin zones from the two sectors are plotted as blue and brown colors, respectively. $G_{1}^{M}$ and $G_{2}^{M}$ are reciprocal lattice vectors.
@@
V.2 Continuum model of twisted trilayer graphene
For TTG, the moiré pattern is formed by the rotation of the top and bottom layer with angle $-\theta/2$ and the middle layer with angle $\theta/2$ relative to $x$ axis respectively. The moiré Brillouin zone is the same as Fig. S3. For TTG, the Hamiltonian reads
$H=H_{t}+H_{m}+H_{b}+H_{int}^{tm}+H_{int}^{bm}.$ (S90)
@@ V.2 Continuum model of twisted trilayer graphene
$H_{t}$ $=\sum_{\xi,\bm{k}}a^{\dagger}_{t,\xi,\bm{k}}\hbar v_{F}\hat{R}_{\theta/2}(\bm{k}-\bm{K_{1}})\cdot(\xi\sigma_{x},\sigma_{y})a_{t,\xi,\bm{k}}~{},$ (S91)
$H_{m}$ $=\sum_{\xi,\bm{k}}a^{\dagger}_{m,\xi,\bm{k}}\hbar v_{F}\hat{R}_{-\theta/2}(\bm{k}-\bm{K_{2}})\cdot(\xi\sigma_{x},\sigma_{y})a_{m,\xi,\bm{k}}~{},$ (S92)
$H_{b}$ $=\sum_{\xi,\bm{k}}a^{\dagger}_{b,\xi,\bm{k}}\hbar v_{F}\hat{R}_{\theta/2}(\bm{k}-\bm{K_{1}})\cdot(\xi\sigma_{x},\sigma_{y})a_{b,\xi,\bm{k}},$ (S93)
@@
$H_{int}^{tm}=\sum_{\xi}a^{\dagger}_{t,\xi,\bm{k}}\left[\left(\begin{array}[]{cc}w_{0}&w_{1}\\ w_{1}&w_{0}\\ \end{array}\right)\delta_{\bm{k},\bm{k}^{\prime}}+\left(\begin{array}[]{cc}w_{0}&w_{1}e^{-i\frac{2\pi}{3}}\\ w_{1}e^{i\frac{2\pi}{3}}&w_{0}\\ \end{array}\right)\delta_{\bm{k+\bm{G}_{1}^{M}},\bm{k}^{\prime}}+\left(\begin{array}[]{cc}w_{0}&w_{1}e^{i\frac{2\pi}{3}}\\ w_{1}e^{-i\frac{2\pi}{3}}&w_{0}\\ \end{array}\right)\delta_{\bm{k}+\bm{G}_{1}^{M}+\bm{G}_{2}^{M},\bm{k}^{\prime}}\right]a_{m,\xi,\bm{k}^{\prime}}+h.c.$ (S94)
$H_{int}^{bm}=\sum_{\xi}a^{\dagger}_{b,\xi,\bm{k}}\left[\left(\begin{array}[]{cc}w_{0}&w_{1}\\ w_{1}&w_{0}\\ \end{array}\right)\delta_{\bm{k},\bm{k}^{\prime}}+\left(\begin{array}[]{cc}w_{0}&w_{1}e^{-i\frac{2\pi}{3}}\\ w_{1}e^{i\frac{2\pi}{3}}&w_{0}\\ \end{array}\right)\delta_{\bm{k+\bm{G}_{1}^{M}},\bm{k}^{\prime}}+\left(\begin{array}[]{cc}w_{0}&w_{1}e^{i\frac{2\pi}{3}}\\ w_{1}e^{-i\frac{2\pi}{3}}&w_{0}\\ \end{array}\right)\delta_{\bm{k}+\bm{G}_{1}^{M}+\bm{G}_{2}^{M},\bm{k}^{\prime}}\right]a_{m,\xi,\bm{k}^{\prime}}+h.c.$ (S95)
@@
V.3 Continuum model of twisted double bilayer graphene
@@
$H_{u}=\sum_{\xi,\bm{k}}a^{\dagger}_{u,\xi,,\bm{k}}\left(\begin{array}[]{cc}H_{1}&U\\ U^{\dagger}&H_{2}\end{array}\right)a_{u,\xi,\bm{k}}~{},$ (S98)
$H_{d}=\sum_{\xi,\bm{k}}a^{\dagger}_{d,\xi,\bm{k}}\left(\begin{array}[]{cc}H_{3}&U\\ U^{\dagger}&H_{4}\end{array}\right)a_{d,\xi,\bm{k}}~{}.$ (S101)
@@
$H_{1}$ $=\hbar v_{0}\hat{R}_{\theta/2}(\bm{k}-\bm{K_{1}})\cdot(\xi\sigma_{x},\sigma_{y})+\frac{V}{2}\sigma_{0}~{},$ (S102)
$H_{2}$ $=\hbar v_{0}\hat{R}_{\theta/2}(\bm{k}-\bm{K_{1}})\cdot(\xi\sigma_{x},\sigma_{y})+\frac{V}{6}\sigma_{0}~{},$ (S103)
$H_{3}$ $=\hbar v_{0}\hat{R}_{-\theta/2}(\bm{k}-\bm{K_{2}})\cdot(\xi\sigma_{x},\sigma_{y})-\frac{V}{6}\sigma_{0}~{},$ (S104)
$H_{4}$ $=\hbar v_{0}\hat{R}_{-\theta/2}(\bm{k}-\bm{K_{2}})\cdot(\xi\sigma_{x},\sigma_{y})-\frac{V}{2}\sigma_{0}~{}.$ (S105)
@@
$U=\left(\begin{array}[]{cc}-v_{4}(\xi k_{x}-ik_{y})&v_{3}(\xi k_{x}+ik_{y})\\ \gamma_{1}&-v_{4}(\xi k_{x}-ik_{y})\\ \end{array}\right)~{}.$ (S106)
@@
$H_{int}=\sum_{\xi}\int_{\bm{r}}\psi^{\dagger}_{u,\xi}(\bm{r})\left(\begin{array}[]{cc}0&0\\ T(\bm{r})&0\\ \end{array}\right)\psi_{d,\xi}(\bm{r})+h.c.~{}.$ (S107)
@@
We adopt the parameters with $(\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4})=(2610,361,283,140)$ meV, and $v_{i}=\frac{\sqrt{3}}{2\hbar}a_{0}\gamma_{i}$ Lee et al. (2019). The interlayer coupling is $(w_{0},w_{1})=(88,100)$ meV. The effect of an out-of-plane electric field is introduced by $V=40$ meV. Once the Bloch wavefunction is obtained, the quantum metric $\mathcal{G}_{ab}$ can be evaluated by Eq. S46. Because of the time-reversal symmetry, we have $\mathcal{G}^{\xi=+}_{ab}(\bm{k})=\mathcal{G}^{\xi=-}_{ab}(-\bm{k})$. For TBG and TTG, since the quantum metric is divergent at the Dirac points, in our calculations we add a small sublattice potential $\Delta_{p}\sigma_{z}$ with $\Delta_{p}=0.01$ to the bottom layer of graphene. This is reasonable because the superconductivity occurs near half filling of the flat band, not near the Dirac points.
Figure S4: (a) The band structure of TBG for $\theta=1.08^{\circ},1.1^{\circ},1.12^{\circ}$. At $\theta=1.08^{\circ}$, the bandwith $W<1$ meV. (b)-(d) The calculation of the superconducting coherence length $\xi$ as a function of $\mu$ for $\theta=1.08^{\circ},1.1^{\circ},1.12^{\circ}$. In (b), $\xi$ from band dispersion $\leq$ 2nm and the quantum metric contribution dominates. When the twist angle is larger, the Fermi velocity changes dramatically, leading to the enhanced contribution from band dispersion in (c) and (d). In (c) and (d), $\xi$ get divergent when $\mu$ is away from half-filling because of the weak pairing gap. The half-filling ($\nu=-0.5$) is labeled by the orange dashed line. The attractive interaction $U_{0}=0.6$ meV.
Figure S5: The self-consistent calculation of the pairing gap $\Delta$ as a function of $\mu$ for different $\theta$. The attractive interaction $U_{0}=0.6$ meV.
V.4 Coherence length of TBG: Mean-field study
@@
$H_{\mathrm{BdG}}=\left(\begin{array}[]{cccc}E_{+\uparrow,\bm{k}}-\mu&0&0&\Delta\\ 0&E_{-\downarrow,\bm{k}}-\mu&-\Delta&0\\ 0&-\Delta&-E_{+\uparrow,-\bm{k}}+\mu&0\\ \Delta&0&0&-E_{-\downarrow,-\bm{k}}+\mu\end{array}\right).$ (S108)
@@
We summarize the calculation of the zero temperature superconducting coherence length $\xi$ in Fig. S5. We calculate the $\xi$ by using Eq.(22) and (23). In our calculation, we use $U_{0}=0.6$ meV, yielding a critical temperature $T_{c}\approx 1.7$K at half filling for $\theta=1.08^{\circ}$. We use the notation $\xi_{bd},\xi_{qm},\xi_{t}$ to denote the coherence length from band dispersion, quantum metric, and total coherence length, respectively.
Near the magic angle ($\theta=1.08^{\circ}$), the bandwidth of TBG changes dramatically as shown in Fig. S4a. At $\theta=1.08^{\circ}$, the density of states near the Dirac point is almost unchanged and $\Delta\approx 0.25$ meV. Near half filling ($\nu=-0.5$), $\xi_{bd}\leq 2$nm due to the narrow band case and $\xi_{qm}$ is about $15$ nm as shown in Fig. S4b. When $\mu$ is away from half filling ($\mu<0.5$meV), $\xi$ from band dispersion is largely enhanced because the pairing is weak in this regime (see Fig. S5a).
When the twist angle is slightly away from $1.08^{\circ}$, the band structure and the Fermi velocity $v_{F}$ change dramatically. As shown in Fig. S4(b)-(d), $\xi_{bd}$ is gradually enhanced. For example, at $\theta=1.1^{\circ}$, near half filling ($\nu=-0.5$), $\xi_{bd}\approx 10$ nm, which is comparable to $\xi_{qm}$. At $\theta=1.12^{\circ}$, $\xi_{bd}$ further get enhanced.

### v6

@@ VIII Competing Interests
$H_{int}=\sum_{\xi}a^{\dagger}_{1,\xi,\bm{k}}[\left(\begin{array}[]{cc}w_{0}&w_{1}\\ w_{1}&w_{0}\\ \end{array}\right)\delta_{\bm{k},\bm{k}^{\prime}}+\left(\begin{array}[]{cc}w_{0}&w_{1}e^{-i\frac{2\pi}{3}}\\ w_{1}e^{i\frac{2\pi}{3}}&w_{0}\\ \end{array}\right)\delta_{\bm{k+\bm{G}_{1}^{M}},\bm{k}^{\prime}}+\left(\begin{array}[]{cc}w_{0}&w_{1}e^{i\frac{2\pi}{3}}\\ w_{1}e^{-i\frac{2\pi}{3}}&w_{0}\\ \end{array}\right)\delta_{\bm{k}+\bm{G}_{1}^{M}+\bm{G}_{2}^{M},\bm{k}^{\prime}}]a_{2,\xi,\bm{k}^{\prime}}+h.c.~{}.$ (91)
Supplementary Fig. 3: The moiré Brillouin zone (mBZ) of moiré graphene family. The original Brillouin zones from the two sectors are plotted as blue and brown colors, respectively. $G_{1}^{M}$ and $G_{2}^{M}$ are reciprocal lattice vectors.
@@
Continuum model of twisted trilayer graphene
For TTG, the moiré pattern is formed by the rotation of the top and bottom layer with angle $-\theta/2$ and the middle layer with angle $\theta/2$ relative to $x$ axis respectively. The moiré Brillouin zone is the same as Supplementary Fig. 2. For TTG, the Hamiltonian reads
$H=H_{t}+H_{m}+H_{b}+H_{int}^{tm}+H_{int}^{bm}.$ (92)
@@
$H_{t}$ $=\sum_{\xi,\bm{k}}a^{\dagger}_{t,\xi,\bm{k}}\hbar v_{F}\hat{R}_{\theta/2}(\bm{k}-\bm{K_{1}})\cdot(\xi\sigma_{x},\sigma_{y})a_{t,\xi,\bm{k}}~{},$ (93)
$H_{m}$ $=\sum_{\xi,\bm{k}}a^{\dagger}_{m,\xi,\bm{k}}\hbar v_{F}\hat{R}_{-\theta/2}(\bm{k}-\bm{K_{2}})\cdot(\xi\sigma_{x},\sigma_{y})a_{m,\xi,\bm{k}}~{},$ (94)
$H_{b}$ $=\sum_{\xi,\bm{k}}a^{\dagger}_{b,\xi,\bm{k}}\hbar v_{F}\hat{R}_{\theta/2}(\bm{k}-\bm{K_{1}})\cdot(\xi\sigma_{x},\sigma_{y})a_{b,\xi,\bm{k}},$ (95)
@@
$H_{int}^{tm}=\sum_{\xi}a^{\dagger}_{t,\xi,\bm{k}}\left[\left(\begin{array}[]{cc}w_{0}&w_{1}\\ w_{1}&w_{0}\\ \end{array}\right)\delta_{\bm{k},\bm{k}^{\prime}}+\left(\begin{array}[]{cc}w_{0}&w_{1}e^{-i\frac{2\pi}{3}}\\ w_{1}e^{i\frac{2\pi}{3}}&w_{0}\\ \end{array}\right)\delta_{\bm{k+\bm{G}_{1}^{M}},\bm{k}^{\prime}}+\left(\begin{array}[]{cc}w_{0}&w_{1}e^{i\frac{2\pi}{3}}\\ w_{1}e^{-i\frac{2\pi}{3}}&w_{0}\\ \end{array}\right)\delta_{\bm{k}+\bm{G}_{1}^{M}+\bm{G}_{2}^{M},\bm{k}^{\prime}}\right]a_{m,\xi,\bm{k}^{\prime}}+h.c.$ (96)
$H_{int}^{bm}=\sum_{\xi}a^{\dagger}_{b,\xi,\bm{k}}\left[\left(\begin{array}[]{cc}w_{0}&w_{1}\\ w_{1}&w_{0}\\ \end{array}\right)\delta_{\bm{k},\bm{k}^{\prime}}+\left(\begin{array}[]{cc}w_{0}&w_{1}e^{-i\frac{2\pi}{3}}\\ w_{1}e^{i\frac{2\pi}{3}}&w_{0}\\ \end{array}\right)\delta_{\bm{k+\bm{G}_{1}^{M}},\bm{k}^{\prime}}+\left(\begin{array}[]{cc}w_{0}&w_{1}e^{i\frac{2\pi}{3}}\\ w_{1}e^{-i\frac{2\pi}{3}}&w_{0}\\ \end{array}\right)\delta_{\bm{k}+\bm{G}_{1}^{M}+\bm{G}_{2}^{M},\bm{k}^{\prime}}\right]a_{m,\xi,\bm{k}^{\prime}}+h.c.$ (97)
@@
Continuum model of twisted double bilayer graphene
@@
$H_{u}=\sum_{\xi,\bm{k}}a^{\dagger}_{u,\xi,,\bm{k}}\left(\begin{array}[]{cc}H_{1}&U\\ U^{\dagger}&H_{2}\end{array}\right)a_{u,\xi,\bm{k}}~{},$ (100)
$H_{d}=\sum_{\xi,\bm{k}}a^{\dagger}_{d,\xi,\bm{k}}\left(\begin{array}[]{cc}H_{3}&U\\ U^{\dagger}&H_{4}\end{array}\right)a_{d,\xi,\bm{k}}~{}.$ (103)
@@
$H_{1}$ $=\hbar v_{0}\hat{R}_{\theta/2}(\bm{k}-\bm{K_{1}})\cdot(\xi\sigma_{x},\sigma_{y})+\frac{V}{2}\sigma_{0}~{},$ (104)
$H_{2}$ $=\hbar v_{0}\hat{R}_{\theta/2}(\bm{k}-\bm{K_{1}})\cdot(\xi\sigma_{x},\sigma_{y})+\frac{V}{6}\sigma_{0}~{},$ (105)
$H_{3}$ $=\hbar v_{0}\hat{R}_{-\theta/2}(\bm{k}-\bm{K_{2}})\cdot(\xi\sigma_{x},\sigma_{y})-\frac{V}{6}\sigma_{0}~{},$ (106)
$H_{4}$ $=\hbar v_{0}\hat{R}_{-\theta/2}(\bm{k}-\bm{K_{2}})\cdot(\xi\sigma_{x},\sigma_{y})-\frac{V}{2}\sigma_{0}~{}.$ (107)
@@
$U=\left(\begin{array}[]{cc}-v_{4}(\xi k_{x}-ik_{y})&v_{3}(\xi k_{x}+ik_{y})\\ \gamma_{1}&-v_{4}(\xi k_{x}-ik_{y})\\ \end{array}\right)~{}.$ (108)
@@
$H_{int}=\sum_{\xi}\int_{\bm{r}}\psi^{\dagger}_{u,\xi}(\bm{r})\left(\begin{array}[]{cc}0&0\\ T(\bm{r})&0\\ \end{array}\right)\psi_{d,\xi}(\bm{r})+h.c.~{}.$ (109)
@@
We adopt the parameters with $(\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4})=(2610,361,283,140)$ meV, and $v_{i}=\frac{\sqrt{3}}{2\hbar}a_{0}\gamma_{i}$ Lee et al. (2019). The interlayer coupling is $(w_{0},w_{1})=(88,100)$ meV. The effect of an out-of-plane electric field is introduced by $V=40$ meV. Once the Bloch wavefunction is obtained, the quantum metric $\mathcal{G}_{ab}$ can be evaluated by Eq. 47. Because of the time-reversal symmetry, we have $\mathcal{G}^{\xi=+}_{ab}(\bm{k})=\mathcal{G}^{\xi=-}_{ab}(-\bm{k})$. For TBG and TTG, since the quantum metric is divergent at the Dirac points, in our calculations we add a small sublattice potential $\Delta_{p}\sigma_{z}$ with $\Delta_{p}=0.01$ to the bottom layer of graphene. This is reasonable because the superconductivity occurs near half filling of the flat band, not near the Dirac points.
Supplementary Fig. 4: (a) The band structure of TBG for $\theta=1.08^{\circ},1.1^{\circ},1.12^{\circ}$. At $\theta=1.08^{\circ}$, the bandwith $W<1$ meV. (b)-(d) The calculation of the superconducting coherence length $\xi$ as a function of $\mu$ for $\theta=1.08^{\circ},1.1^{\circ},1.12^{\circ}$. In (b), $\xi$ from band dispersion $\leq$ 2nm and the quantum metric contribution dominates. When the twist angle is larger, the Fermi velocity changes dramatically, leading to the enhanced contribution from band dispersion in (c) and (d). In (c) and (d), $\xi$ get divergent when $\mu$ is away from half-filling because of the weak pairing gap. The half-filling ($\nu=-0.5$) is labeled by the orange dashed line. The attractive interaction $U_{0}=0.6$ meV.
Supplementary Fig. 5: The self-consistent calculation of the pairing gap $\Delta$ as a function of $\mu$ for different $\theta$. The attractive interaction $U_{0}=0.6$ meV.
Coherence length of TBG: Mean-field study
@@
$H_{\mathrm{BdG}}=\left(\begin{array}[]{cccc}E_{+\uparrow,\bm{k}}-\mu&0&0&\Delta\\ 0&E_{-\downarrow,\bm{k}}-\mu&-\Delta&0\\ 0&-\Delta&-E_{+\uparrow,-\bm{k}}+\mu&0\\ \Delta&0&0&-E_{-\downarrow,-\bm{k}}+\mu\end{array}\right).$ (110)
@@
We summarize the calculation of the zero temperature superconducting coherence length $\xi$ in Supplementary Fig. 4. We calculate the $\xi$ by using Eq.(22) and (23). In our calculation, we use $U_{0}=0.6$ meV, yielding a critical temperature $T_{c}\approx 1.7$K at half filling for $\theta=1.08^{\circ}$. We use the notation $\xi_{bd},\xi_{qm},\xi_{t}$ to denote the coherence length from band dispersion, quantum metric, and total coherence length, respectively.
Near the magic angle ($\theta=1.08^{\circ}$), the bandwidth of TBG changes dramatically as shown in Supplementary Fig. 4a. At $\theta=1.08^{\circ}$, the density of states near the Dirac point is almost unchanged and $\Delta\approx 0.25$ meV. Near half filling ($\nu=-0.5$), $\xi_{bd}\leq 2$nm due to the narrow band case and $\xi_{qm}$ is about $15$ nm as shown in Supplementary Fig. 4b. When $\mu$ is away from half filling ($\mu<0.5$meV), $\xi$ from band dispersion is largely enhanced because the pairing is weak in this regime (see Supplementary Fig. 5a).
When the twist angle is slightly away from $1.08^{\circ}$, the band structure and the Fermi velocity $v_{F}$ change dramatically. As shown in Supplementary Fig. 4(b)-(d), $\xi_{bd}$ is gradually enhanced. For example, at $\theta=1.1^{\circ}$, near half filling ($\nu=-0.5$), $\xi_{bd}\approx 10$ nm, which is comparable to $\xi_{qm}$. At $\theta=1.12^{\circ}$, $\xi_{bd}$ further get enhanced.
Supplementary References
  • Huhtinen et al. (2022) Kukka-Emilia Huhtinen, Jonah Herzog-Arbeitman, Aaron Chew, Bogdan A. Bernevig,  and Päivi Törmä, “Revisiting flat band superconductivity: Dependence on minimal quantum metric and band touchings,” Phys. Rev. B 106, 014518 (2022)arXiv:2203.11133 [cond-mat.supr-con] .
  • Törmä et al. (2018) Päivi Törmä, Long Liang,  and Sebastiano Peotta, “Quantum metric and effective mass of a two-body bound state in a flat band,” Physical Review B 98, 220511 (2018).
  • Hofmann et al. (2023) Johannes S Hofmann, Erez Berg,  and Debanjan Chowdhury, “Superconductivity, charge density wave, and supersolidity in flat bands with a tunable quantum metric,” Physical review letters 130, 226001 (2023).
  • Koshino et al. (2018) Mikito Koshino, Noah F. Q. Yuan, Takashi Koretsune, Masayuki Ochi, Kazuhiko Kuroki,  and Liang Fu, “Maximally Localized Wannier Orbitals and the Extended Hubbard Model for Twisted Bilayer Graphene,” Physical Review X 8, 031087 (2018)arXiv:1805.06819 [cond-mat.mes-hall] .
  • Lee et al. (2019) Jong Yeon Lee, Eslam Khalaf, Shang Liu, Xiaomeng Liu, Zeyu Hao, Philip Kim,  and Ashvin Vishwanath, “Theory of correlated insulating behaviour and spin-triplet superconductivity in twisted double bilayer graphene,” Nature Communications 10, 5333 (2019)arXiv:1903.08685 [cond-mat.str-el] .